OpenStudy (anonymous):

Use the concept of slope to find t such that the three points are collinear: (-6,1), (1,t), (10,5)

OpenStudy (lgbasallote):

the equation for slope is \[\huge \frac{y_2- y_1}{x_2 - x_1}\] if they are all colinear that means their slopes are the same so\[\huge \frac{t-1}{1 - (-6)} = \frac{5-t}{10-1} = \frac{5-1}{10 - (-6)}\] does that help?

OpenStudy (lgbasallote):

@EnergizedStudent15 you there?

OpenStudy (anonymous):

but yes this really does help! Alot! The only thing I'm confused about is answering it :/

OpenStudy (lgbasallote):

here's what you do... \[\huge \frac{t-1}{1-(-6)} = \frac{5-1}{10-(-6)}\] simplify those two sides..you should get an "algebra" situation

OpenStudy (anonymous):

Yea......... ;^-^

OpenStudy (anonymous):

so from there it's pretty much cross multiplication right?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

but why is the \[5-t/10-1\] crossed out?

OpenStudy (lgbasallote):

well you can use that too...it's just that an equation only consists of two parts..left side and right side...so i picked two from the three and equated them

OpenStudy (anonymous):

ohhhh okay.